# Differential torsion sensor for direct detection of ultralight vector dark matter

### v2

@@ III Differential torsion sensor
Assuming 5-kg test masses on both ends of each 0.6-m sensor beam, we have a total moment of inertia of 0.75 kg m2 for each pendulum, equivalent to having a $\sim 5$ kg sensor beam with additional mounting equipment and a mechanical lever arm from the axis of rotation to the center of each test mass of 0.24 m. The torsion pendulums (suspended by two 0.6-m long tungsten suspension wires with a $Q$ factor of $3\times 10^{3}$ Cavalleri et al. (2009)) have a mechanical resonant frequency of $0.026$ Hz. The common suspension point of the two pendulums provides a large suppression on mechanical common modes. With a four-stage mechanical suspension and isolation chain ($1.5$ m in diameter and $3$ m tall) operating in a vacuum tank, a suspension point motion of $10^{-11}$ m Hz-1/2 and $10^{-11}$ rad Hz-1/2 at $0.1$ Hz can be achieved Chua et al. (2023). This isolation performance guarantees a suspension thermal noise limited differential angular sensitivity of $6.7\times 10^{-13}$ rad Hz-1/2 at 0.1 Hz at a temperature of 300 K. We consider an impedance-matched cavity with finesse of $\sim 150$. With an incident laser power of 10 mW at a wavelength of $1.064\,\mu{\rm m}$ and using the optimal modulation depth, the displacement-equivalent shot noise in the cavities is $4\times 10^{-18}$ m Hz-1/2. With a 0.24-m mechanical lever arm, the shot noise limited differential angular sensitivity is at the level of $10^{-17}$ rad Hz-1/2. Both the beam vibration thermal noise and the mirror thermal noise are subdominant. Using beam tubes made of Niobium (with the first vibration mode frequency at $\sim 480$ Hz and a $Q$ factor of $10^{6}$ Duffy and Umstattd (1994)) and with a laser spot size of 1-mm diameter on the fused silica cavity mirrors, both the beam vibration thermal noise and the mirror substrate and coating thermal noise levels Braginsky and Vyatchanin (2003); Levin (1998) are below the suspension thermal noise at lower frequencies and shot noise at higher frequencies. With this setup, the projected amplitude spectrum density of the differential torque at design sensitivity, denoted by $S_{\delta\tau}^{1/2}$, is presented in Figure 2 Chua et al. (2023).
@@ IV Sensitivity
A more accurate sensitivity estimate is considered for future work, which should include as-built experimental implementations, software simulations, and rigorous analysis methods. To achieve the desired sensitivity, other sources of noise apart from thermal noise and shot noise, e.g., Newtonian noise, magnetic noise, etc., need to be well monitored and controlled. Magnetic noise can be mitigated by placing the experiment in a magnetically shielded room or placing the torsion beams in an enclosed shield made from mu-metal. In addition, potential eddy currents are minimized by using low magnetic permeability materials near the torsion pendulums. Newtonian noise, which cannot be shielded or attenuated, requires different techniques from mechanical isolation to be mitigated. In the most sensitive frequency band of $\sim 0.3$ Hz, the Newtonian noise can reach an order of magnitude higher than the total noise budget shown in Figure 2 Chua et al. (2023). There are multiple pathways to reduce the Newtonian noise to a subdominant level. An optimal solution is to build a second, closely positioned, identical differential torsion sensor equipped with the four test masses made of the same material, which will measure the Newtonian noise, allowing it to be canceled in the dark matter sensor. Other mitigation options include: 1) Placing the experiment underground can mitigate the impact of surface waves and body waves to some extent Punturo et al. (2010). 2) Excavating earth beneath the vacuum tank and replacing it with a lightweight fill can suppress local density and displacement fluctuations Hall et al. (2021). 3) A seismometer array can be effective in recording the seismic field in the vicinity of the differential torsion sensor and estimating the Newtonian noise contribution Hall et al. (2021).
@@
Figure 3: Theoretically projected searched sensitivity on the $B\!-\!L$ coupling constant (black band), assuming $T_{\rm obs}=180$ day. The lower and upper bounds correspond to setting a coherent integration time using Equations (15) and (16), respectively. The existing limits set by the Eöt-Wash experiment Schlamminger et al. (2008b), MICROSCOPE Bergé et al. (2018); Touboul et al. (2017), and the recent single torsion balance experiment by Shaw et al. Shaw et al. (2022), are plotted for comparison.

### v3

@@ III Differential torsion sensor
Assuming 5-kg test masses on both ends of each 0.6-m sensor beam, we have a total moment of inertia of 0.75 kg m2 for each pendulum, equivalent to having a $\sim 5$ kg sensor beam with additional mounting equipment and a mechanical lever arm from the axis of rotation to the center of each test mass of 0.24 m. The torsion pendulums (suspended by two 0.6-m long tungsten suspension wires with a $Q$ factor of $3\times 10^{3}$ Cavalleri et al. (2009)) have a mechanical resonant frequency of $0.026$ Hz. The design incorporates a common suspension point of the two pendulums and provides a large suppression on mechanical common modes. The four-stage mechanical suspension and isolation chain ($1.5$ m in diameter and $3$ m tall), including low-noise sensors (such as Nanometrics Trillium 240 Lantz and Kissel (2023)) and active feedback and controls, operating in a vacuum tank, can achieve a suspension point motion of $10^{-11}$ m Hz-1/2 and $10^{-11}$ rad Hz-1/2 at $0.1$ Hz Chua et al. (2023). This isolation performance guarantees a suspension thermal noise limited differential angular sensitivity of $6.7\times 10^{-13}$ rad Hz-1/2 at 0.1 Hz at a temperature of 300 K. We consider an impedance-matched cavity with finesse of $\sim 150$. With an incident laser power of 10 mW at a wavelength of $1.064\,\mu{\rm m}$ and using the optimal modulation depth, the displacement-equivalent shot noise in the cavities is $4\times 10^{-18}$ m Hz-1/2. With a 0.24-m mechanical lever arm, the shot noise limited differential angular sensitivity is at the level of $10^{-17}$ rad Hz-1/2. Key formulae for suspension thermal noise and shot noise are provided in Appendix A. Both the beam vibration thermal noise and the mirror thermal noise are subdominant. Using beam tubes made of Niobium (with the first vibration mode frequency at $\sim 480$ Hz and a $Q$ factor of $10^{6}$ Duffy and Umstattd (1994)) and with a laser spot size of 1-mm diameter on the fused silica cavity mirrors, both the beam vibration thermal noise and the mirror substrate and coating thermal noise levels Braginsky and Vyatchanin (2003); Levin (1998) are below the suspension thermal noise at lower frequencies and shot noise at higher frequencies. With this setup, the projected amplitude spectrum density of the differential torque at design sensitivity, denoted by $S_{\delta\tau}^{1/2}$, is presented in Figure 2 Chua et al. (2023).
@@ IV Sensitivity
A more accurate sensitivity estimate is considered for future work, which should include as-built experimental implementations, software simulations, and rigorous analysis methods. To achieve the desired sensitivity, other sources of noise apart from thermal noise and shot noise, e.g., Newtonian noise, magnetic noise, etc., need to be well monitored and controlled. To ensure the magnetic noise is at least one order of magnitude below the noise budget, we require a field $\lesssim 100\mu T$ (c.f. the environment background is at the level of $\sim 50\mu T$). Magnetic noise can be further mitigated by placing the torsion beams in a cubic-meter-scaled enclosed shield made from mu-metal if necessary. In addition, potential eddy currents are minimized by using low magnetic permeability materials close to the torsion pendulums within the mu-metal enclosure. Newtonian noise, which cannot be shielded or attenuated, requires different techniques from mechanical isolation to be mitigated. We need one order of magnitude Newtonian noise mitigation with the experiment located in a low-seismic-noise environment equivalent to Peterson’s new low-noise model  Peterson (1993). There are multiple pathways to reduce the Newtonian noise to a subdominant level. An optimal solution is to build a second, closely positioned, identical differential torsion sensor equipped with the four test masses made of the same material, which will measure the Newtonian noise, allowing it to be canceled in the dark matter sensor. Other mitigation options include: 1) Placing the experiment underground can mitigate the impact of surface waves and body waves to some extent Punturo et al. (2010). 2) Excavating earth beneath the vacuum tank and replacing it with a lightweight fill can suppress local density and displacement fluctuations Hall et al. (2021). 3) A seismometer array can be effective in recording the seismic field in the vicinity of the differential torsion sensor and estimating the Newtonian noise contribution Hall et al. (2021).
@@
Figure 3: Theoretically projected searched sensitivity on the $B\!-\!L$ coupling constant (black band), assuming $T_{\rm obs}=180$ day. The lower and upper bounds correspond to setting a coherent integration time using Equations (15) and (16), respectively. The existing limits set by the Eöt-Wash experiment Schlamminger et al. (2008b), MICROSCOPE Bergé et al. (2018); Touboul et al. (2017), the recent single torsion balance experiment by Shaw et al. Shaw et al. (2022), and LISA Pathfinder Frerick et al. (2024), are plotted for comparison.
@@ Acknowledgements.
Appendix A Useful formulae for noise estimates
The power spectral density of suspension thermal noise (in units of m2 Hz-1), as a function of angular frequency $\omega$, is derived from the fluctuation dissipation theorem Callen and Welton (1951); Callen and Greene (1952) and is given by Saulson (1990); Cumming et al. (2009, 2012)
$S_{x}(\omega)=\frac{4k_{B}T}{\omega^{2}}\Re(Y_{x}),$ (17)
where $k_{B}$ is Boltzmann’s constant, $T$ is the absolute temperature, and $Y_{x}$ is the complex mechanical admittance of the torsion pendulum.
The shot-noise-equivalent displacement of a cavity is given by Black (2001)
$S_{\rm shot}=\frac{1}{8\mathcal{F}}\left(\frac{hc\lambda}{P_{c}}\right)^{1/2},$ (18)
where $\mathcal{F}$ is the finesse of the optical cavity, $h$ is Planck’s constant, $c$ is the speed of light, $\lambda$ is the laser wavelength, and $P_{c}$ is the power in the carrier signal.